# JudgeAgent: Dynamically Evaluate LLMs with Agent-as-Interviewer

### v1

@@ 3.3 Evaluation Feedback
This process constitutes the result–generation phase of JudgeAgent. Following the Benchmark Grading and multiple rounds of Interactive Extension, JudgeAgent aggregates all question–answer performances of the target model across varying difficulty. JudgeAgent then evaluates the target from multiple dimensions, including potential deficiencies in semantic understanding, logical reasoning, and detail capture, as well as its comprehensive performance, producing a final evaluation report in textual form. In addition, JudgeAgent simultaneously provides feedback and suggestions aimed at addressing the identified capability gaps, serving as a reference for model optimization.
Inspired by Generative Reward Model methods(Ankner et al., 2024; Ye et al., 2024; Li et al., 2025), we propose a streamlined procedure for verifying the validity of capability evaluations. For each batch of questions, JudgeAgent generates corresponding feedback and suggestions. Using these suggestions as guidance, the target model is prompted to answer the same set of questions again. By comparing accuracy before and after the intervention, we can indirectly validate the effectiveness of the evaluation to some extent.
@@ 4 Experiments
In this section, we validate the effectiveness of the evaluation results from JudgeAgent using the method introduced in Section 3.3. The detailed process is illustrated in Algorithm 2. The following research questions guide our experiments: First, does JudgeAgent genuinely discover the shortcomings of the target model (RQ1)? Second, to what extent does each evaluation mechanism in JudgeAgent influence the evaluation results (RQ2)?
Experiment Setup. We select GPT-4.1111We use gpt-4.1-2025-04-14 from OpenAI’s official API for all experiments(Achiam et al., 2023) to be the core LLM in JudgeAgent, which serves as both a generator and evaluator model. The Interactive Extension in our experiments is limited to a maximum of 3 rounds.
@@ 4.1 Dataset and Target Model Selection
The dynamic evaluation of JudgeAgent begins with Benchmark Grading, a phase that requires a predefined static dataset. In our experiments, we utilized datasets including: MedQA(Jin et al., 2021), MultiHop-RAG(Tang & Yang, 2024), and QuALITY(Pang et al., 2022).
MedQA contains multiple-choice questions in the style of the Medical Licensing Examination questions, and we select the US subset of MedQA, which has questions in English. MultiHop-RAG consists of phrase Q&A queries constructed from news articles published between September and December 2023. We use questions without reference knowledge from MedQA and MultiHop-RAG to evaluate the target model’s knowledge proficiency and to verify whether the JudgeAgent can discover the target’s knowledge gaps through dynamic evaluation. QuALITY is a multiple-choice question dataset for long document comprehension. The questions are written and validated by contributors who have read the entire passage. When employing QuALITY questions to test the target model, background passages are provided, as QuALITY is designed to assess comprehension and inference based on these passages, which primarily consist of fictional narratives. Therefore, unlike MedQA and MultiHop-RAG, we do not use QuALITY to validate the JudgeAgent’s effectiveness in evaluating knowledge gaps, but rather its ability to guide comprehension and reasoning. We select the validation set of QuALITY to conduct experiments, since the test set doesn’t provide correct answers. Although QuALITY has already categorized difficulty into 0 and 1 based on whether the annotator’s accuracy under the speed validation setting exceeds 50%, we divided the original question difficulty into three tiers of Hard, Medium, and Easy, using 20% and 60% as the thresholds to better align with the JudgeAgent setup.
@@ 4.2 Evaluation Metrics
We utilize Accuracy to measure the performance of the target model on base datasets. To validate the effectiveness of the evaluation suggestions from JudgeAgent, we propose two metrics, Correction Rate and Correct-to-Error Rate. The detailed definition of these metrics is as follows: (1) Accuracy (ACC). This is a common metric to measure the performance, which is the proportion of questions correctly answered by the target model. (2) Correction Rate (CR). This metric quantifies the proportion of questions that the target model initially answered incorrectly but subsequently answered correctly after receiving evaluation suggestions, which measures the effectiveness of the evaluation suggestions. A higher correction rate indicates better performance of the evaluation suggestions. (3) Correct-to-Error Rate (CtE). This metric serves as the inverse of the Correction Rate, representing the proportion of questions that were initially answered correctly but subsequently answered incorrectly. A lower CtE indicates greater effectiveness of the evaluation suggestions.
@@ 4.3 Main Results and Analysis
Table 1: The results on MedQA and MultiHopRAG, and all values are percentages. ACC1 and ACC2 refer to the accuracy of the target model before and after receiving evaluation suggestions.
@@
Based on the performance across different difficulty levels of QuALITY as shown in Table 2, we can analyze the effectiveness of JudgeAgent’s evaluation suggestions for different target models on questions of various difficulties. For relatively stronger models (Qwen3, GPT-4.1, and Gemini-2.5-pro), JudgeAgent provides stronger guidance and optimization, particularly on questions of higher difficulty (Medium and Hard). In contrast, for the relatively weaker models, JudgeAgent leads to considerable improvement across all difficulty levels, with particularly notable gains on Easy-level questions compared to stronger models. These results further indicate that JudgeAgent more accurately assesses the capability boundaries of target models and provides more difficulty-adaptive optimization to the target models. Furthermore, for all target models, JudgeAgent’s evaluation suggestions demonstrate clear optimizing guidance on Hard-level questions that involve complex reasoning, suggesting that JudgeAgent effectively identifies underlying deficiencies in target models through dynamic interactive evaluation.
@@ 4.4 Ablation Study
To address RQ2, we conduct ablation studies on MedQA with GLM4-Flash as the target model. Under the condition of maintaining the target model’s responses to base questions unchanged, we removed different key modules and conducted the evaluation pipeline of JudgeAgent. By analyzing and comparing the performance improvement of the target model under various ablation settings, we investigated the extent to which different mechanisms influence the evaluation results. The results of the ablation study are shown in Table 3.
Table 3: The results of the ablation study with MedQA as the base dataset, and all values are percentages. $\Delta=\text{ACC1}-\text{ACC2}=\text{CR}-\text{CtE}$ refers to the overall improvement on the target models after receiving evaluation from JudgeAgent.
@@
   -w/o interacitve extension 93.01 1.97 0.67 1.30
JudgeAgent (w/o context graph) removes the context graph and only uses chunks sampled randomly from the original knowledge base to generate questions. The context graph ensures knowledge relevance between newly synthesized questions and the base questions. The results indicate that removing the knowledge relevance provided by the context graph reduces the improvement effect of the JudgeAgent’s evaluation feedback on the target model. A possible explanation is that the context graph provides relevance between the extended questions and the base question in terms of focused entities and knowledge, thereby allowing evaluation feedback about the extended questions to supplement the background knowledge of the base question. In contrast, under this ablation setting, the lack of knowledge relevance leads to more fragmented evaluation suggestions from JudgeAgent, which may instead disrupt the thinking of target models. The higher correction-to-error rate (CtE) compared to the setting -w/o difficulty-adaptive also provides supporting evidence for this potential interference.
JudgeAgent (w/o difficulty-adaptive) removes the difficulty-adaptive mechanism and generates questions with fixed difficulty rules in the prompt. The difficulty-adaptive mechanism enables the JudgeAgent to assess the proficiency level of the target model more precisely. The results show that after removing the difficulty-adaptive mechanism, the effectiveness of JudgeAgent’s evaluation feedback was somewhat diminished. This suggests that without the difficulty-adaptive mechanism, JudgeAgent may provide evaluation feedback that deviates from the target model’s actual proficiency level, thereby compromising the effectiveness of the evaluation results.
JudgeAgent (w/o interactive extension) removes the Interactive Extension phase and only evaluates the target models with base datasets. The Interactive Extension serves as the core module of the JudgeAgent’s dynamic evaluation process, expanding both the scope of knowledge and the range of capabilities involved in the evaluation. The results show that removing the Interactive Extension significantly weakens the evaluative effectiveness of JudgeAgent. Compared to other ablation settings, a clear decrease in CR and an increase in CtE can be observed. This indicates that the comprehensive evaluation enabled by the dynamic expansion of questioning, which covers both breadth of knowledge and depth of capabilities, is crucial to the final evaluation results. Evaluating solely based on the base questions may prevent the evaluator from fully gauging the target model’s proficiency level. It is on this foundation that further improvements, such as those brought by knowledge relevance and difficulty adaptation, are built.
@@ 4.5 Case Study
To gain an in-depth understanding of JudgeAgent’s evaluation mechanism and its effectiveness, we analyze a comparative case of GLM4-Flash’s responses to MedQA questions by receiving evaluation suggestions from direct evaluation and JudgeAgent. A brief overview of the comparative case is shown in Figure 2, while the detailed content is provided in Appendix D.
Figure 2: The brief overview of the comparative case of LLM’s performance by receiving suggestions from direct evaluation and JudgeAgent.
In this case, the target LLM provided an incorrect answer when queried on a base question. When directly evaluated by an LLM evaluator without being supplied with the correct answer or relevant background knowledge, the LLM evaluator adopted a more conservative approach, in which the LLM evaluator opted to generically enumerate the potential consequences of closed abdomnial injury across various anatomical sites, rather than performing targeted differentiation between the specific medical concepts involved. Consequently, even after receiving the evaluator’s feedback, the target model continued to produce an incorrect answer. This persistence is attributable to the fact that peritonitis symptoms resulting from both colon injury and duodenal bulb injury can exhibit delayed onset, a nuance the model failed to discern.
In contrast, JudgeAgent first extracted two key entities from the base question, closed abdomnial injury and peritonitis. It then retrieved relevant knowledge entities and text chunks, such as splenic rupture and duodenal injury, from the context graph. Based on the sampled knowledge paths, JudgeAgent generated a series of extended questions. By analyzing the target model’s performance on these extended questions, JudgeAgent identified that the target model had insufficient understanding of the symptoms associated with duodenal injury caused by closed abdominal injury, thereby providing targeted knowledge guidance. Ultimately, with this feedback, the target model successfully answered the original base question correctly.
The comparative case above demonstrates that through dynamic evaluation paradigm based on knowledge-driven synthesis and difficulty-adaptive mechanism, JudgeAgent is capable of accurately identifying gaps in the target model’s knowledge and providing targeted evaluation feedback.
@@ 5 Conclusion
In this paper, we propose JudgeAgent, a knowledge-target adaptive dynamic evaluation framework based on a new interviewer-style evaluation paradigm. JudgeAgent precisely assesses the knowledge and capabilities boundaries and deficiencies of target models and provides targeted evaluation suggestions by a comprehensive evaluation pipeline, which consists of benchmark grading, interactive extension, and evaluation feedback. We also introduce a novel insight into validating the effectiveness of evaluation results. Our experiments based on these validation methods demonstrate the effectiveness of JudgeAgent and its dynamic evaluation paradigm. In our future work, we will further refine the novel evaluation paradigms and dynamic evaluation frameworks and develop more reliable evaluation tools for LLMs.
Appendix A Construction of Context Graph

### v2

@@ 3.3 Evaluation Feedback
This process constitutes the result generation phase. JudgeAgent aggregates the target’s all Q&A performances in the above two stages, and evaluates the target from multiple dimensions, including knowledge boundaries, logical reasoning, and comprehensive performance, producing a text evaluation report. Based on this evaluation, JudgeAgent simultaneously provides suggestions aimed at addressing the identified capability gaps, serving as a reference for model optimization.
Inspired by Generative Reward Model methods (Ankner et al., 2024; Ye et al., 2024; Li et al., 2025) and the evaluation report method proposed by (Kim et al., 2025), JudgeAgent generates feedback and suggestions based on all Q&A history for each batch, without directly providing the correct answer or background knowledge. Benefiting from context graphs, the suggestions can identify key knowledge concepts and extend relevant information to provide concise and valuable feedback, rather than providing ”cheating information” specific to seed questions. To verify the suggestion’s effectiveness, the target model is prompted to answer the same questions again with these suggestions. By comparing accuracy before and after the intervention, we can indirectly validate the effectiveness of the evaluations.
@@ 4 Experiments
In this section, we validate the effectiveness of JudgeAgent’s evaluations using the method in Section 3.3, which is detailed in Algorithm 2. The following research questions guide our experiments: First, does JudgeAgent genuinely discover the shortcomings of the target model (RQ1)? Second, to what extent does each mechanism in JudgeAgent influence the evaluations (RQ2)?
Experiment Setup. We select GPT-4.1111We use gpt-4.1-2025-04-14 from OpenAI’s official API for all experiments(Achiam et al., 2023) to be the core LLM in JudgeAgent, which serves as both a generator and evaluator model. In our experiments, the batch size in Benchmark Grading is 3, and the Interactive Extension is limited to a maximum of 3 rounds.
@@ 4.1 Dataset and Target Model Selection
The initial stage of JudgeAgent, Benchmark Grading, requires a predefined static benchmark. In our experiments, we select MedQA(Jin et al., 2021), MultiHop-RAG(Tang & Yang, 2024), and QuALITY(Pang et al., 2022). We remove the background knowledge of questions from MedQA and MultiHop-RAG during evaluation to evaluate the target’s knowledge boundaries rather than its comprehension ability, and to verify whether JudgeAgent can discover the target’s knowledge deficiencies. We use QuALITY to validate the effectiveness of JudgeAgent in guiding the comprehension and reasoning of target models. For detailed statistical information and the preprocessing procedures of these datasets in our experiments, refer to Appendix D.
@@ 4.2 Evaluation Metrics
The metrics used in our experiments are as follows:
(1) Accuracy (ACC). We use this metric to measure the performance of the target model on base datasets, which is the proportion of questions correctly answered by the target model. To investigate the validity of JudgeAgent’s evaluations, we compared the changes in the ACC before and after dynamic evaluation. ACC1 represents the target model’s ACC in answering base questions in Benchmark Grading stage, measuring its performance before receiving evaluation feedback. ACC2 denotes the ACC in answering the same questions after receiving feedback from JudgeAgent.
(2) Correction Rate (CR). This metric quantifies the proportion of questions that the target model initially answered incorrectly but subsequently answered correctly after receiving evaluation suggestions, which measures the effectiveness of the feedback suggestions from JudgeAgent. A higher correction rate indicates better performance of the suggestions.
(3) Correct-to-Error Rate (CtE). This metric serves as the inverse of the Correction Rate, representing the proportion of questions that were initially answered correctly but subsequently answered incorrectly. A lower CtE indicates greater effectiveness of the suggestions.
@@ 4.3 Main Results and Analysis
Table 1: The results on MedQA and MultiHopRAG, and all values are percentages.
@@
Based on the performance across different difficulties of QuALITY as shown in Table 2, we can analyze the effectiveness of JudgeAgent’s suggestions for different targets on various difficulties. For stronger models (Qwen3, GPT-4.1, and Gemini-2.5-pro), JudgeAgent provides stronger guidance and optimization, particularly on questions of higher difficulty (Medium and Hard), since these models may have already mastered the basic knowledge of Easy-level questions. In contrast, for the weaker models, JudgeAgent leads to considerable improvement across all difficulty levels, with particularly notable gains on the Easy level, since JudgeAgent’s feedback is more effective at filling gaps in basic concepts. These results further indicate that JudgeAgent more accurately assesses the capability boundaries of targets and provides more difficulty-adaptive optimization to the targets. Furthermore, for all target models, JudgeAgent’s suggestions demonstrate clear optimizing guidance on Hard-level questions that involve complex reasoning, suggesting that JudgeAgent effectively identifies underlying deficiencies in target models through dynamic interactive evaluation.
@@
In addition, we conducted cross-validation experiments to verify that the JudgeAgent’s suggestions are effective not only for base questions, further corroborating the validity of the main experiment results. The results and analysis are detailed in Appendix E.2
@@ 4.4 Ablation Study
To address RQ2, we conduct ablation studies on MedQA with GLM4-Flash as the target LLM. When maintaining the responses to base questions unchanged, we removed different modules of JudgeAgnet. By comparing the performance improvement under various ablation settings, we investigated how much different modules influence the evaluations. The results are shown in Table 3.
Table 3: The results of the ablation study with MedQA as the base dataset, and all values are percentages. $\Delta=\text{ACC2}-\text{ACC1}=\text{CR}-\text{CtE}$ refers to the overall improvement on the target models after receiving evaluation from JudgeAgent.
@@
   -w/o interactive extension 93.01 1.97 0.67 1.30
JudgeAgent (w/o context graph) removes the context graph, which ensures the knowledge relevance between synthesized questions and the base ones, and only uses chunks sampled randomly from the original knowledge base. The results indicate that removing the context graph reduces the improvement effect of the JudgeAgent’s evaluations. A possible explanation is that the lack of context graph prevents extended questions and base questions from being associated in terms of focused entities and knowledge, leading to fragmented suggestions without accurate information, which may disrupt the thinking of target LLMs. The higher correction-to-error rate (CtE) compared to the setting -w/o difficulty-adaptive also provides supporting evidence for this potential interference.
JudgeAgent (w/o difficulty-adaptive) removes the difficulty-adaptive mechanism, which enables JudgeAgent to assess the target model’s capability precisely, and generates questions with fixed difficulty rules in the prompt. The results show that the removal of this module diminishes the effectiveness of JudgeAgent’s evaluations, demonstrating the importance of the difficulty-adaptive mechanism for providing effective evaluations.
JudgeAgent (w/o interactive extension) removes the Interactive Extension, which is the core process of dynamic evaluation to extend the evaluation scope of knowledge and capability, and only evaluates models with base datasets. The results show that removing the Interactive Extension significantly weakens the effectiveness of JudgeAgent compared to other ablation settings. This indicates that the evaluation enabled by the dynamic expansion, which expands both breadth and depth of knowledge and capability in dynamic evaluation, is crucial to the final evaluation suggestions.
@@
4.5 Parameter Analysis
Figure 3: The results of different expansion rounds on MedQA and MultiHop-RAG. @K indicates the ACC improvement after receiving suggestions from the K-th interaction.
Are more expansion rounds better for JudgeAgent? To answer this research question, we tested the evaluation effectiveness of JudgeAgent under different rounds in the Interactive Extension, as shown in Figure 3. The results demonstrate that as the expansion rounds increase, the effectiveness of JudgeAgent, which is represented by the accuracy improvement of target models, also gradually improves. However, the trend gradually slows down, showing a relatively clear marginal effect.
As the rounds increase, JudgeAgent can expand more questions around the knowledge related to base questions. After reaching certain rounds, the Q&A history is sufficient for JudgeAgent to identify the target’s knowledge deficiencies around base questions. Additional questions serve only as corroborative rather than critical evidence. There is a clear marginal effect in Figure 3. For example, the results of GPT-4.1 on MultiHop-RAG, the improvement per round decreases gradually from 2.04% by the first round, 1.13% by the second, 0.71% by the third, and to 0.23% by the last round, and other curves exhibit similar patterns. Therefore, JudgeAgent expands a maximum of 3 rounds in our experiments to avoid wasting resources and time.
@@ 4.6 Case Study
To gain a deeper understanding of Agent-as-Interviewer’s mechanism, we analyze GLM4-Flash’s responses to MedQA questions by receiving suggestions from direct evaluation and Agent-as-Interviewer. The seed questions where the target answers correctly are omitted in this case. A brief overview of the case is shown in Figure 4, while the detailed content is provided in Appendix H.
Figure 4: The brief overview of the comparative case in the Case Study.
In this case, the target LLM answers the base question incorrectly. When directly evaluated by the LLM evaluator, the evaluator adopted a more conservative approach to generically enumerate the potential consequences of closed abdominal injury, rather than to perform a comparison between the specific medical concepts involved. Consequently, even after receiving the feedback, the target LLM still outputs an incorrect answer, which is attributable to a nuance it failed to discern: peritonitis symptoms resulting from both colon injury and duodenal bulb injury can exhibit delayed onset.
In contrast, JudgeAgent first extracted two key entities, closed abdominal injury and peritonitis. It then retrieved relevant entities with texts, such as splenic rupture and duodenal injury, from the context graph. Based on the sampled knowledge paths, JudgeAgent generated a series of extended questions. Given that the target model consistently errs on all questions where duodenal injury and closed abdominal injury co-occur, JudgeAgent can determine that the target model had insufficient understanding of this concept entity, thereby providing specific knowledge guidance. Ultimately, the target LLM successfully answered the original base question correctly with this feedback.
The case above demonstrates that through the Agent-as-Interviewer paradigm, JudgeAgent is capable of accurately identifying gaps in the target’s knowledge and providing targeted feedback.
@@ 5 Conclusion
In this paper, we propose Agent-as-Interviewer, a dynamic evaluation paradigm that employs LLM agents to conduct multi-turn interactions for evaluation. This paradigm utilizes agents to call knowledge tools for wider and deeper knowledge in the dynamic question generation to achieve more complete evaluations, and leverages agents to plan query strategies for adjustment of the question difficulty to match the actual capabilities of target LLMs. Based on this paradigm, we develop JudgeAgent, a knowledge-wise dynamic evaluation framework to assess LLMs in multi-turn Q&A sessions and provide valuable suggestions that assist target LLMs in optimizing themselves. Thorough experiments validate that Agent-as-Interviewer can precisely assess the target’s knowledge boundaries. In our future work, we will further refine this novel evaluation paradigm and develop more reliable evaluation frameworks for LLMs.
Appendix A The Use of Large Language Models
In this paper, we only used LLMs, including DeepSeek-R1 and ChatGPT, for polishing the writing. Specifically, we used LLMs to assist in refining the language, improving readability, and grammar checking. The authors take full responsibility for the content of the paper.
Appendix B Analysis and Proofs of Adaptive Difficulty-Control Mechanisms
In Section 3.1, we have formalized the difficulty control mechanisms as follows:
Let $d_{1},d_{2},d_{3}]$ respectively denote the difficulty Easy, Medium, and Hard, $n_{it},n_{if}$ respectively represent the number of correct and incorrect answers for $d_{i}$-level questions, $c_{it}$ be the score gain for a correct answer at $d_{i}$, and $t_{ij}$ be the score threshold between $d_{i}$ and $d_{j}$, we can get the average score $avg_{s}$ and the difficulty of the next round question generation $d_{next}$:
[…]
where $c_{1t},c_{2t},c_{3t}]=[1,1.5,2]$ and $t_{12},t_{23}]=[0.5,1]$ in our experiments. In this section, we will illustrate the criteria and rules of the score gain $c_{it}$ and the threshold $t_{ij}$, along with the analysis and proofs. For ease of representation, let $N=\sum_{i=1}^{3}(n_{it}+n_{if})$ in the subsequent analysis.
For the selection of score gains and thresholds, we followed simple cognitive principles to establish several rules that a reasonable difficulty control mechanism should satisfy:
R1. Balance: Correctly answering a more difficult question should gain a higher score. If the number of correctly answered Easy and Hard questions is equal, the level should be considered equivalent to the same total number of Medium questions. Additionally, the conventional range of the average score should correspond equally to the three difficulty levels in ascending order.
R2. Generalizability: The same formula, score gains, and thresholds should apply even to questions without difficulty labels, where difficulty is estimated solely based on the accuracy.
R3. Improvability: Correctly answering questions should increase the average score, and there should be a possibility of exceeding the threshold to advance to the next difficulty level.
R4. Stability: If the capability estimation remains at a certain difficulty for a long period, the likelihood of advancing to the next level should gradually decrease.
Following the above rules, we can analyze the mathematical conditions that $c_{it}$ and $t_{ij}$ must satisfy. From R1, we can get that $c_{1t}+c_{3t}=2c_{2t}\land c_{1t}<c_{2t}<c_{3t}$. Since we have set the score for correctly answering a question without a difficulty level as $c_{2t}$, the range of $avg_{s}$ can be determined as $0\,,\,c_{2t}]$, representing the spectrum from answering all questions incorrectly to answering all correctly. Based on the condition about threshold in R1, we can get that $t_{12}=c_{2t}/3\ ,\ t_{23}=2c_{2t}/3$ in the case that difficulty levels are absent. Due to R2, these threshold conditions can be generalized to the case where questions are provided with difficulty levels.
For R3, let $avg_{s}$ denote the average score after answering $N$ questions, and $avg_{s}^{\prime}$ denote the average score after answering $N+1$ questions. Then R3 is equivalent to satisfying the following conditions:
$\textbf{C1.}\quad avg_{s}^{\prime}=\frac{avg_{s}N+c_{it}}{N+1}>avg_{s},\quad\text{holds true always when}\ \ avg_{s}\leq t_{(i,i+1)}$ (2)
$\textbf{C2.}\quad\exists\ avg_{s0},N_{0},\quad\text{s.t.}\ avg_{s}^{\prime}=\frac{avg_{s0}N_{0}+c_{it}}{N_{0}+1}>t_{(i,i+1)}$
Simplifying Eq 2.C1, we find that $c_{it}>avg_{s}$ holds true always when $avg_{s}\leq t_{(i,i+1)}$. Therefore, we can conclude that $c_{it}>t_{(i,i+1)}$, meaning $c_{2t}>c_{1t}>t_{12}$ and $t_{23}<c_{2t}<c_{3t}$. Simplifying Eq 2.C2, we can obtain a relationship between $avg_{s0}$ and $N_{0}$:
$avg_{s0}>t_{(i,i+1)}-\frac{c_{it}-t_{(i,i+1)}}{N_{0}}$ (3)
[…]
Finally, R4 can be stated as follows: given the average score $a\leq t_{(i,i+1)}$ after $N$ questions, the number $n$ of consecutive questions that must be answered correctly to surpass the level threshold should increase with $N$. We can obtain the following inequality:
$\frac{aN+c_{it}n}{N+n}>t_{(i,i+1)}\Rightarrow(c_{it}-t_{(i,i+1)})n>(t_{i,i+1}-a)N$ (4)
[…]
$c_{1t}<c_{2t}<c_{3t}$ (5)
$c_{1t}+c_{3t}=2c_{2t}$
$t_{12}=c_{2t}/3\ ,\ t_{23}=2c_{2t}/3$
$c_{it}>t_{(i,i+1)}\Rightarrow c_{2t}>c_{1t}>t_{12}\ ,\ t_{23}<c_{2t}<c_{3t}$
Therefor, we first set $c_{2t}=1.5$, and so $t_{12}=0.5\ ,\ t_{23}=1$. Then, since $1.5=c_{2t}>c_{1t}>t_{12}=0.5$ and $1=t_{23}<c_{2t}=1.5<c_{3t}$, we set $c_{1t}=1$ for convenience, and so $c_{3t}=2c_{2t}-c_{1t}=2$.
Appendix C Construction of Context Graph